OpenStudy (anonymous):

Solve: (2x + 5)(x - 3) = 6

OpenStudy (campbell_st):

expand 2x^2 -x - 15 = 6 2x^2 - x - 21 = 0 (2x -7)(x +3) = 0 x = 7/2 or -3